A 120-V hair dryer has two settings: 850 W and 1550 W .

B) Determine the resistance at 850 W setting.

C) Determine the resistance at 1550 W setting.
Answer #1

power P=\frac{W}{t}=\frac{vq}{t}

P=vi \left (i=\frac{q}{t} \right )

P=\frac{v^{2}}{R}\left (\therefore i=\frac{v}{R} \right )

we know v=120 V

b)resistance of hair dryer at p=850 W

R=\frac{v^{2}}{P}

R=\frac{120^{2}}{850}

=16.94\Omega

c)resistance of hair dryer at 1550 W

R=\frac{120^{2}}{1550}

=9.29 \Omega